# Atomic Calibration of LLMs in Long-Form Generations

### v2

@@ 1 Introduction
While large language models (LLMs) (Touvron et al., 2023; Jiang et al., 2023; OpenAI, 2022) excel in various tasks, they still struggle with trustworthiness issues. LLMs often suffer from hallucinations, generating factually inaccurate content and misleading responses (Zhang et al., 2023b; Huang et al., 2023), which limits their application in high-risk real-world scenarios (Hu et al., 2023). To address this, confidence calibration aims to estimate the underlying uncertainty of model predictions and reflect the true likelihood of correctness (Guo et al., 2017). A calibrated model is crucial for real-world applications, as it allows us to determine the extent to which we can trust models’ predictions (Zhu et al., 2023; Mahaut et al., 2024). Improved calibration enables more reliable confidence estimation, warning users when not to trust the model and thus mitigating the impact of hallucinations.
Most existing work on LLM calibration focuses on short-form QA tasks (Jiang et al., 2021; Tian et al., 2023; Zhu et al., 2023; Ulmer et al., 2024), using datasets like TriviaQA and Natural Questions (Joshi et al., 2017), where answers are typically under 10 words. In contrast, real-world queries often elicit much longer responses (Zhang et al., 2024; Yang et al., 2024b), spanning hundreds or thousands of words. In such cases, response quality is not simply binary, as answers may mix accurate and inaccurate statements.
Recent work has begun to address calibration in long-form generation (Zhang et al., 2024; Huang et al., 2024; Liu et al., 2023; Fadeeva et al., 2024; Jiang et al., 2024). Several approaches estimate a single confidence score for the entire response (macro calibration; upper, Figure 1), while others assess confidence at the level of atomic claims (Liu et al., 2023; Fadeeva et al., 2024; Jiang et al., 2024) (atomic calibration; lower, Figure 1). However, previous work leaves several key research questions unanswered: Why is it important to evaluate calibration at the atomic-claim level? What factors influence calibration results at this level? What patterns can be observed by analyzing calibration at the atomic-claim level?
@@ 2 Related Work
Uncertainty and Calibration in Long-form Generations. Existing research on uncertainty estimation and calibration primarily focuses on multiple-choice or short-form questions (Zhu et al., 2023; Kuhn et al., 2022; Lin et al., 2023; Tian et al., 2023; Ulmer et al., 2024). There is an increasing interest on calibration for long-form generations. Huang et al. (2024) proposed a unified calibration framework for all text generation tasks, comparing distributions of both correctness and the associated confidence of responses. Band et al. (2024) introduced linguistic calibration, where models explicitly express their uncertainty during long-form generation. Zhang et al. (2024) proposed Luq, an uncertainty estimation method tailored to long-form generation, demonstrating its effectiveness in ensembling different LLMs. Another line of work (Liu et al., 2023; Fadeeva et al., 2024; Jiang et al., 2024; Yuan et al., 2024) decomposes sentences into atomic claims and assigns confidence scores to each claim. However, a unified definition of atomic-level calibration remains lacking. Clarifying this concept and identifying key influencing factors are essential steps toward improving calibration in long-form generation.
@@ 3 Atomic Calibration
⇏$\not\Rightarrow$ $\frac{1}{N}\sum_{i=1}^{N}\mathbb{P}\left(y_{i}\mid f(c_{i})_{y_{i}}=\beta_{i}\right)=\beta$
⇏$\not\Rightarrow$ $\mathbb{P}\left(y_{i}\mid f(c_{i})_{y_{i}}=\beta_{i}\right)=\beta,\forall i\in\{1,....,N\}.$
@@ 4.1 Generative Methods
Generative methods assume that the consistency between different generation samples provides a reliable estimation of model uncertainty (Zhang et al., 2024; Jiang et al., 2024). Generally, an additional natural language inference (NLI) model is used to calculate the consistency. In particular, we have the following two variations:
@@ 4.3 Confidence Fusion Strategies
Combining confidence scores has proven effective for calibration (Huang et al., 2024; Rivera et al., 2024), but existing methods typically only use a single fixed weight, $\alpha$, to combine the scores, ignoring the confidence disagreement. For instance, Rivera et al. (2024) computes the weighted average (WAvg) for confidence scores $A$ and $B$: $C=A\cdot\alpha+B\cdot(1-\alpha)$. This approach does not account for the agreement between the two scores. For example, when $\alpha=0.5$, confidences of 0 and 1 are treated the same as confidences of 0.4 and 0.6, although the former may indicate higher uncertainty due to a larger disagreement. To address this, we propose two simple but effective methods that consider confidence disagreement $d=B-A$.
@@ 5.1 Experiment Setup
Models. We utilize seven LLMs from three model families with varying sizes: Llama3 Instruct (8B and 70B) (Meta, 2024), Mistral Instruct (7B and 8x7B) (Jiang et al., 2023), and Qwen2 Instruct (7B, 52B-A14B, and 72B) (Yang et al., 2024a).
@@ 6.3 The Utilities of Atomic Calibration
While the primary goal of atomic calibration is to provide fine-grained calibration evaluation for models, we also explore its utilities in several downstream tasks, including: (1) Selective Question Answering (Kamath et al., 2020; Cole et al., 2023; Yang et al., 2023), which involves setting a confidence threshold to selectively reject low-confidence answers, ensuring that only high-confidence responses are retained; (2) LLM-Ensemble (Zhang et al., 2024), which leverages multiple models to generate responses to the same question, selecting the answer with the highest confidence, thereby combining the strengths of each model; and (3) Atomic Claims Reunion (Thirukovalluru et al., 2024; Jiang et al., 2024), which involves sampling multiple responses, breaking then into atomic claims, evaluating their confidence, and reassembling only high-confidence claims to produce a more reliable final answer. Among these applications, we observe consistent improvements in factuality with atomic-level examination. Detailed experimental settings and results can be found in Appendix C.
It is important to note that, unlike previous work on Selective Question Answering (Huang et al., 2024) and LLM-Ensemble (Zhang et al., 2024) for long-form generation, which mainly rely on atomic-level confidence estimation to enhance the overall quality of responses (with responses either being entirely accepted or rejected), Atomic Claims Reunion does not require an overall response-level score. Instead, it relies entirely on the confidence of atomic claims to select and combine the most accurate claims. This means the final answer may contain claims from different sampled answers. More importantly, we observe that models with better atomic-level calibration (e.g., Qwen2 in Table 7, Appendix C) exhibit greater improvements after the reunion process, emphasizing the importance of examining and refining atomic calibration.
@@ Appendix A Atomic Calibration Metrics
Following Zhang et al. (2024), we calculate Spearman Correlation to assess whether samples with higher factuality have correspondingly higher confidence scores. Compared to Pearson Correlation, it focuses on assessing the rank correlation, is robust to outliers and does not require that data is in normal distribution.
@@ Appendix B Statistics in Atomic and Macro Calibration
Figure 7: Answer with less than 10 atomic facts (%) by dataset and model. Notably, Mistral-8$\times$7B-Instruct has 6% short answers. Human evaluation reveals that these responses are primarily instances where the model refuses to answer.

### v3

@@ 1 Introduction
While large language models (LLMs) (Touvron et al., 2023; Jiang et al., 2023; OpenAI, 2022) excel in various tasks, they still struggle with trustworthiness issues. LLMs often suffer from hallucinations, generating factually inaccurate content and misleading responses (Zhang et al., 2024a; Shelmanov et al., 2025; He et al., 2025), which limits their application in high-risk real-world scenarios (Hu et al., 2023). To address this, confidence calibration aims to estimate the underlying uncertainty of model predictions and reflect the true likelihood of correctness (Guo et al., 2017; Zhang et al., 2025a, b). In this work, we distinguish between calibration as a property and as a process. We use the term atomic calibration to refer to the property of each atomic claim’s confidence being well-aligned with its factuality (Zhou et al., 2025). In contrast, the extra post-hoc methods used to achieve this alignment (such as temperature scaling and platt scaling) are referred to as calibration methods or a calibrator, which is not considered in this paper.
A calibrated model is crucial for real-world applications, as it allows us to determine the extent to which we can trust models’ predictions (Zhu et al., 2023; Mahaut et al., 2024). Improved calibration enables more reliable confidence estimation, warning users when not to trust the model and thus mitigating the impact of hallucinations.
Most existing work on LLM calibration focuses on short-form QA tasks (Jiang et al., 2021; Tian et al., 2023; Zhu et al., 2023; Ulmer et al., 2024), using datasets like TriviaQA and Natural Questions (Joshi et al., 2017), where answers are typically under 10 words. In contrast, real-world queries often elicit much longer responses (Zhang et al., 2024b; Yang et al., 2025a, b), spanning hundreds or thousands of words. In such cases, response quality is not simply binary, as answers may mix accurate and inaccurate statements.
Recent work has begun to address calibration in long-form generation (Zhang et al., 2024b; Huang et al., 2024; Liu et al., 2023; Fadeeva et al., 2024; Jiang et al., 2024). Several approaches estimate a single confidence score for the entire response (macro calibration; upper, Figure 1), while others assess confidence at the level of atomic claims (Liu et al., 2023; Fadeeva et al., 2024; Jiang et al., 2024) (atomic calibration; lower, Figure 1). However, previous work leaves several key research questions unanswered: Why is it important to evaluate calibration at the atomic-claim level? What factors influence calibration results at this level? What patterns can be observed by analyzing calibration at the atomic-claim level?
@@ 2 Related Work
Uncertainty and Calibration in Long-form Generations. Existing research on uncertainty estimation and calibration primarily focuses on multiple-choice or short-form questions (Zhu et al., 2023; Kuhn et al., 2022; Lin et al., 2023; Tian et al., 2023; Ulmer et al., 2024). There is an increasing interest on calibration for long-form generations. Huang et al. (2024) proposed a unified calibration framework for all text generation tasks, comparing distributions of both correctness and the associated confidence of responses. Band et al. (2024) introduced linguistic calibration, where models explicitly express their uncertainty during long-form generation. Zhang et al. (2024b) proposed Luq, an uncertainty estimation method tailored to long-form generation, demonstrating its effectiveness in ensembling different LLMs. Another line of work (Liu et al., 2023; Fadeeva et al., 2024; Jiang et al., 2024; Yuan et al., 2024) decomposes sentences into atomic claims and assigns confidence scores to each claim. However, a unified definition of atomic-level calibration remains lacking. Clarifying this concept and identifying key influencing factors are essential steps toward improving calibration in long-form generation.
@@ 3 Atomic Calibration
$\not\Rightarrow$ $\frac{1}{N}\sum_{i=1}^{N}\mathbb{P}\left(y_{i}\mid f(c_{i})_{y_{i}}=\beta_{i}\right)=\beta$
$\not\Rightarrow$ $\mathbb{P}\left(y_{i}\mid f(c_{i})_{y_{i}}=\beta_{i}\right)=\beta,\forall i\in\{1,..,N\}.$
@@ 4.1 Generative Methods
Generative methods assume that the consistency between different generation samples provides a reliable estimation of model uncertainty (Zhang et al., 2024b; Jiang et al., 2024). Generally, an additional natural language inference (NLI) model is used to calculate the consistency. In particular, we have the following two variations:
@@ 4.3 Confidence Fusion Strategies
Combining confidence scores has proven effective for calibration (Huang et al., 2024; Rivera et al., 2024), but existing methods typically only use a single fixed weight, $\alpha$, to combine the scores, ignoring the confidence disagreement. For instance, Rivera et al. (2024) computes the weighted average (WAvg) for confidence scores $A$ and $B$: $C=A\cdot\alpha+B\cdot(1-\alpha)$. This approach does not account for the agreement between the two scores. For example, when $\alpha=0.5$, confidences of 0 and 1 are treated the same as confidences of 0.4 and 0.6, although the former may indicate higher uncertainty due to a larger disagreement. To address this, we propose two simple but effective methods that consider confidence disagreement d=BAd=B-A.
@@ 5.1 Experiment Setup
Models. We utilize seven LLMs from three model families with varying sizes: Llama3 Instruct (8B and 70B) (Meta, 2024), Mistral Instruct (7B and 8x7B) (Jiang et al., 2023), and Qwen2 Instruct (7B, 52B-A14B, and 72B) (Yang et al., 2024).
@@ 6.3 The Utilities of Atomic Calibration
While the primary goal of atomic calibration is to provide fine-grained calibration evaluation for models, we also explore its utilities in several downstream tasks, including: (1) Selective Question Answering (Kamath et al., 2020; Cole et al., 2023; Yang et al., 2023), which involves setting a confidence threshold to selectively reject low-confidence answers, ensuring that only high-confidence responses are retained; (2) LLM-Ensemble (Zhang et al., 2024b), which leverages multiple models to generate responses to the same question, selecting the answer with the highest confidence, thereby combining the strengths of each model; and (3) Atomic Claims Reunion (Thirukovalluru et al., 2024; Jiang et al., 2024), which involves sampling multiple responses, breaking then into atomic claims, evaluating their confidence, and reassembling only high-confidence claims to produce a more reliable final answer. Among these applications, we observe consistent improvements in factuality with atomic-level examination. Detailed experimental settings and results can be found in Appendix C.
It is important to note that, unlike previous work on Selective Question Answering (Huang et al., 2024) and LLM-Ensemble (Zhang et al., 2024b) for long-form generation, which mainly rely on atomic-level confidence estimation to enhance the overall quality of responses (with responses either being entirely accepted or rejected), Atomic Claims Reunion does not require an overall response-level score. Instead, it relies entirely on the confidence of atomic claims to select and combine the most accurate claims. This means the final answer may contain claims from different sampled answers. More importantly, we observe that models with better atomic-level calibration (e.g., Qwen2 in Table 7, Appendix C) exhibit greater improvements after the reunion process, emphasizing the importance of examining and refining atomic calibration.
@@ Appendix A Atomic Calibration Metrics
Following Zhang et al. (2024b), we calculate Spearman Correlation to assess whether samples with higher factuality have correspondingly higher confidence scores. Compared to Pearson Correlation, it focuses on assessing the rank correlation, is robust to outliers and does not require that data is in normal distribution.
@@ Appendix B Statistics in Atomic and Macro Calibration
[…]